OpenStudy (anonymous):

find the volume of a basketball in cubic inches if the circumference of a basketball is 30 inches

OpenStudy (anonymous):

\[4\pi*r ^{3}\]

OpenStudy (anonymous):

just use that formula and you'll get the volume

OpenStudy (anx):

find the radius, C= pi(d) so 30(pi) = d = 9.5 d= 9.5 / 2 = 4.8 v = 4/3 (pi)r^3 so 4/3(pi)(4.8)^3 = 463.25 in^3

OpenStudy (anx):

the basketball is a sphere so use volume of a sphere

OpenStudy (anonymous):

V 4 r3 3 4 9.55 3 V3 2 V 456.05 in3

OpenStudy (anonymous):

\[v \approx456.05inches ^{3}\]